# Strain stiffening due to stretching of entangled particles in random packings of granular materials

### v1

@@ IV.3.1 Probability of stretching and local deformation ratio
Figure 13: Probability $p_{s}$ of staples being plastically deformed so as to indicate tensional loads on particles, as a function of maximum packing strain $\gamma_{max}$. At small strain $\gamma_{max}$ the typical loads may be too small to cause permanent plastic deformation. Line: proportional fit for $\gamma_{max}>0.2$.
@@
The probability of stretching $p_{s}$ is measured as the fraction of staples in the categories with extended arms (with and without backbone bending) based on the categorization in Fig. 12. This probability $p_{s}$ is plotted as a function of the maximum strain $\gamma_{max}$ of each experiment in Fig. 13, with error bars representing (Poisson) counting statistics. For small $\gamma_{max}\leq 0.15$, we find almost zero deformed particles. In this range, it is likely that most bent particles had forces on them less than the yield force or elastic deformation limit (about 9 N from Fig. 24 in Appendix G), and so do not produce permanent deformation. Thus, we are unable to infer stress distributions from plastic deformation of particles in this range. For larger $\gamma_{max}>0.2$, we find significant fractions of bent particles. To obtain a model proportionality for $p_{s}$, we fit a proportionality for $\gamma_{max}\geq 0.2$, resulting in $p_{s}=(1.13\pm 0.07)\gamma$. This may underestimate the number of stretched particles if there are a significant number with forces below the elastic deformation limit. For the sake of applying a lowest-order model, we will assume that the linear proportionality of $p_{s}(\gamma)$ extends to the small-$\gamma$ range to account for particles whose forces were below the elastic deformation limit.
@@ IV.3.2 Test of the stress-strain relation
For comparison, a quadratic fit of the stress-strain curve for staples yields $\sigma=(62.76\pm 0.06)\gamma^{2}$ MPa with a standard deviation of 0.5 MPa around the fit. The model curvature is consistent with the fit value within the 37% model uncertainty. Thus, we can attribute the stress-strain relation of packings of staples to the deformation of particles due to tensile loadings.
For completeness, we tested a quadratic fit with two fit parameters which yields $a_{1}=(2.71\pm 0.04)$ MPa and $a_{2}=(59.4\pm 0.2)$ MPa. In this case, the coefficient $a_{2}$ only changes by a few percent, and $a_{1}$ only contributes to $9\%$ of the total stress at $\gamma=0.5$. Thus, the strength of packings of entangling staples can also be approximated by a dominant quadratric function without a linear term, which is again attributed to the linear probability of entanglement $p_{s}$ with increasing strain $\gamma$, similar to chains. With the linear term, the data fits with a reduced $\chi^{2}=1.0$ when using the standard deviation of multiple runs as an uncertainty, indicating the linear term is reproducible in the case of staples, if relatively small.
@@ VI Discussion
It is interesting to consider if this understanding of entanglement could lead to mixtures with enhanced strength or strain stiffening in soils, concrete, and other granular materials. Fibers are typically only added to soils at a fraction less than about 1% to provide reinforcement, and while this reinforcement increases strength, it does not produce strain-stiffening when the fibers are mixed in randomly et al (2022). The small fraction of fibers is likely too little to produce strong strain stiffening from system-filling clusters of entanglements. The same could be said about concrete reinforced with fibers Ashour et al. (1992). Larger packing fractions of fibers have been confirmed to enhance the weaker self-amplified friction mechanism for strain stiffening from localized strain Dumont et al. (2018).
@@ VII Conclusions
Using x-ray tomography measurements of the packing structure, we showed that the number of entanglements $n$ of chains is proportional to an effective free enclosed area $A_{eff}$ surrounded by a chain available to entangle other particles, accounting for excluded volume effects (Fig. 6). This area $A_{eff}$ and the expected number of entanglements $n$ can be calculated given the shape of the particles in the packing assuming random particle arrangements (Fig. 5). The threshold of about 1 entanglement per enclosure required to produce system-filling clusters and strong strain stiffening is similar to the Erdös-Rényi model for random graphs.
We presented this model in a form that could be extended to packings of different particle shapes that can entangle. In particular, for a particular rigid particle geometry, one could calculate from a Monte Carlo simulation the possible entanglement arrangements the effective free enclosed area $A_{eff}$ (Appendix E), which leads to the number of entanglements $n$, and thus whether or not there are enough entanglements to connect in system-filling clusters to produce strong strain stiffening. While these quantities can be predicted from the known geometry of rigid particles (Fig. 15), calculating these quantities required detailed structural information for flexible chains to calculate $A_{eff}$ and $n$.
@@
We presented a model for the stress-strain relation of random packings with system-filling entanglements, where the stress originates from the stress required to deform the particles, which is due to tensional loads on the particles in the cases tested. Using a mean-field model, the slope of the stress-strain curve of the packing was written as a product of the effective particle modulus, the packing fraction, the probability of stretched particles $p_{s}$, and the ratio of particle strain to packing strain (Eq. 2). For both chains and staples, the increase in effective stiffness with strain that characterizes strain stiffening can be attributed to the increase in the fraction of stretched particles $p_{s}(\gamma)$ as packing strain $\gamma$ increases (Fig. 7b, 13) and causes more entangled particles to pull on each other, while the effective particle stiffness and the ratio of particle strain to packing strain are approximately constant (Figs. 17, 24, 14). A roughly linear increase in the probability of stretched particles $p_{s}$ with packing strain $\gamma$ leads to a roughly quadratic stress-strain relation (Fig. 3, 11). The curvature of the stress-strain curve was predicted within the model uncertainties ($\leq 37\%$) of measured values for packings of both chains and staples (Figs. 3, 11). To predict the stress-strain relation, the effective particle modulus and packing fraction are relatively easy to measure, however, the probability of stretching and the ratio of particle strain to packing strain required information from experiments with three-dimensional imaging, permanent plastic deformation, or simulations of packings, and are not yet predictable from first principles.
@@ VIII Acknowledgements
We thank Mike Dunlap for assistance with materials testing equipment, Peter Eng and Mark Rivers for their assistance with x-ray tomography at the Advanced Photon Source, Ling-Nan Zhou for his particle tracking code, Nicholas Rodenberg and Dylan Murphy for preliminary measurements, and Sulimon Sattari for performing simulations of network models for comparison, Henos Musie for performing measurements for Fig. 24, Tammer Abiyu for performing measurements for Fig. 11, David Brantley for performing measurements for Fig. 18, and Xuan-Truc Nguyen for analyzing particles for Figs. 12 and 14. We acknowledge the Advanced Photon Source general user program at Argonne National Laboratory for providing beam time. This work was supported by the NSF MRSEC program under DMR-0820054 and DMR-2011854.
@@ Appendix A Self-amplified friction model comparison
Why is the scaling of our stress-strain curves so different from those of Dumont et al. (2018)? An important difference is that Dumont et al. (2018) used a narrow indenter with an open top surface that allows dilation of the packing without resistance, which results in a more localized strain so a few chains near the indenter are responsible for most of the displacement. We confirmed experimentally that using a narrow indenter and free boundary allows beads to move past each other when the packing is indented. This more localized motion allows the mechanism of self-amplified friction to determine which chains move and thus the response measured by the indenter. In contrast, our experiments used a flat plate that produced a relatively uniform strain in the packing, and a weak confinement from the sides with the elastic membrane. In this case, the entanglements of long chains prevent particles from shearing past each other, thus preventing the self-amplified friction mechanism, and instead requiring load to be generated deformation of particles. This comparison confirms the importance of preventing shear banding Regev and Reichhardt (2013) in the strong strain stiffening behavior that we observed. However, it also highlights that the apparent measured stress-strain relation is not necessarily an intrinsic material property, and even different mechanisms can exist for strain stiffening for the same material depending on whether the boundary conditions allowa more localized shear banding or prevents it.
@@ Appendix C Bend angles of chains
To test whether the compression of tight loops of chains contributes to strong strain stiffening, the bending stiffness of a chain loop was measured by pulling on a looped chain. A chain with $\xi=8$ beads was bent into a loop (shown in the inset of Fig. 18) and held in place between two sheets of acrylic separated by 4.5 mm in such a manner that the chain would not unloop or catch on another bead as the ends of the chain were pulled. Pictures were taken to measure local bend angles $\theta$ between the links of neighboring chains (shown in the inset of Fig. 18) with an uncertainty of 1 degree limited by the image resolution. Four angles were measured for each chain (those centered on the beads with solid circles overlaid in their centers in the inset of Fig. 18), and the measurement was repeated for two different chains to obtain a total of eight measured angles at several different extensional force values. The average angles for each force are shown in Fig. 18, with the error bars indicating the standard deviations of the measured angles. We measured no resolvable force until the chains were bent beyond $45\pm 1^{\circ}$ degrees, which corresponds to the minimum loop circumference $\xi=360^{\circ}/45^{\circ}=8$. While there is an initial linear force $F_{\theta}$ with bend angle $\theta$, above 2.5 N, the chains started to bend non-uniformly, indicated by the large increase in standard deviation of $\theta$ with increasing force shown in Fig. 18; in particular one link would bend much more than the others, indicating yielding of the weakest link.
@@ Appendix D Packing fraction
The global packing fraction was also obtained from the x-ray tomography data at strains $\gamma=0$ and $0.2$ by dividing the sphere volume by the average volume of convex hulls surrounding the beads. The $N=5$ and $N=20$ samples dilated ($\phi$ decreased) under compression, while the $N=10$ sample was initially loose (presumably due to an inconsistency in packing procedure) so compacted ($\phi$ increased) under compression. The packing fraction at strains of $\gamma=0.2$ appears to follow a smooth trend with $N$, which is consistent with the expectation that packings evolve towards a critical packing fraction (a.k.a. critical state) under shear, regardless of initial packing fraction (Scholfield and Wroth, 1968). The initially loose packing for $N=10$ may have contributed to there being very few links at full extension at $\gamma=0$ (Fig. 7b), which made it impossible to resolve the average link strain $\gamma_{p}$ for $N=10$, and about half the number of entanglements as other datasets at $\gamma=0$ (Fig. 6). Once the sample was strained to $\gamma=0.2$, these statistics matched more closely with the data with the other strain stiffening packing at $N=20$.
@@ Appendix E Calculation of effective free enclosed area $A_{eff}$
The ratio of $A_{eff}/A$ is plotted as a function of chain segment length $m$ in Fig. 23. We find a consistent $A_{eff}/A=0.27\pm 0.01$ in the range of strong strain stiffening ($m\geq 8$), with the $\pm$ corresponding to the standard deviation at different contour lengths $m$ for both $N=10$ and 20. This consistency of the ratio $A_{eff}/A$ in this range is in agreement with the earlier expectation that this ratio is a constant for staples Gravish et al. (2012) ($A_{eff}/A$ corresponds to $1/\alpha$ in Gravish et al. (2012)). In the limit of large $m$, the distance from where the entangled chain crosses the entangling manifold to the entangling particle $l(\Phi,x)$ is likely to be large compared to the particle diameter $a$. In that limit, the factor $1-a^{2}/l(\Phi,x)^{2}$ in the integrand of Eq. 10 approaches 1. For chains, this corresponds to the limit of large $m$ which can have entanglement manifolds much wider than $a$. However, for smaller $m\leq 8$, the ratio $A_{eff}/A$ does drop to zero as the constraints of excluded volume become more significant where the entangled width of the entanglement manifold is only slightly larger than $a$ (i.e. the largest $l(\Phi,x)$ is only slightly larger than $a$). Using the derivation of Eq. 10 as a guide, it can be inferred that the maximum possible value of $A_{eff}/A$ is 1/2 in the extreme limit of thin entangled particles inside a large entangled volume where $a/l(\Phi,x)$ would be small for most of the volume.
@@ E.1 Monte Carlo calculations of $A_{eff}/A$
The calculation of $A_{eff}/A$ for chains shows that the number of entanglements $n$ can be predicted given a particle geometry, but since chains have a flexible shape it required input of x-ray tomography data to get the particle geometry in the packing. To confirm that such a geometric calculation for $A_{eff}/A$ can make a priori predictions for rigid particles based on the particle shape without packing information, we carried out a Monte Carlo simulations of $A_{eff}/A$ for staples at various aspect ratios of arm length over backbone length. The geometry of staples is different than Fig. 22, assuming cylindrical arms of diameter $a$ in the enclosed rectangular area $A=(L_{x}-a/2)(L_{y}-a/2)$. We randomly sample values of arms crossing the entangled plane at positions $20\leq x\leq L_{x}-a/$ 2, $20\leq y\leq L_{y}/2-a/$ 2, and orientations $20\leq\Theta\leq\pi/$ 2 (with probability weighted by $\sin\Theta$), and $20\leq\Phi\leq\pi/$ 2. Valid entanglements with no overlap of particles in the set of coordinates are weighted by a factor of $\cos\Theta$ in the average to account for the larger volume of the staple in the entangling volume, and contribute a factor of zero if there is overlap. A set of positions and orientations is considered to be a valid entangled position without overlapping another particle if
@@
The values of $A_{eff}/A$ obtained from this Monte Carlo calculation for staples are shown in Fig. 23b for different aspect ratios $L_{x}/L_{y}$ with a fixed backbone length $L_{y}=13a$, corresponding to the simulation of Gravish et al. (2012).
For the Monte Carlo simulation of $\langle A(\Theta,\Phi)\cos\Theta/r\rangle$ to predict $p_{s}$ for our staples in Sec. V, we include a weighting of $\cos\Theta/r$ instead of $\cos\Theta$ above, and approximate our rectangular cross-section staples by setting the diameter $a=w=1.24$ mm, and use our $L_{x}=4.2$ mm and $L_{y}=11.3$ mm. Due to the weak aspect ratio dependence when $L_{x}\gg a$, the approximation of rectangular arms as circular cross-section cylinders results in no more than a 2% error on $\langle A(\Theta,\Phi)\cos\theta/r\rangle$.

### v2

@@ IV.3.1 Probability of stretching and local deformation ratio
Figure 13: Probability $p_{s}$ of staples being plastically deformed so as to indicate tensional loads on particles, as a function of maximum packing strain $\gamma_{max}$. At small strain $\gamma_{max}$ the typical loads may be too small to cause permanent plastic deformation. The increasing trend $p_{s}(\gamma_{max})$ is predicted by the model to result in strain stiffening. Line: proportional fit for $\gamma_{max}>0.2$.
@@
The probability of stretching $p_{s}$ is measured as the fraction of staples in the categories with extended arms (with and without backbone bending) based on the categorization in Fig. 12. This probability $p_{s}$ is plotted as a function of the maximum strain $\gamma_{max}$ of each experiment in Fig. 13, with error bars representing (Poisson) counting statistics. For small $\gamma_{max}\leq 0.15$, we find almost zero deformed particles. In this range, it is likely that most bent particles had forces on them less than the yield force or elastic deformation limit (about 9 N from Fig. 24 in Appendix G), and so do not produce permanent deformation. Thus, we are unable to infer stress distributions from plastic deformation of particles in this range. For larger $\gamma_{max}>0.2$ mm, we find significant fractions of bent particles. To obtain a coefficient for the model trend $p_{s}(\gamma)$ assuming it is linear, we fit a proportionality coefficient for $\gamma_{max}\geq 0.2$, resulting in $p_{s}=(1.13\pm 0.07)\gamma$. This may underestimate the number of stretched particles if there are a significant number with forces below the elastic deformation limit. For the sake of applying a lowest-order model, we will assume that the linear proportionality of $p_{s}(\gamma)$ extends to the small-$\gamma$ range to account for particles whose forces were below the elastic deformation limit.
@@ IV.3.2 Test of the stress-strain relation
For comparison, a quadratic fit of the stress-strain curve for staples yields $\sigma=(62.76\pm 0.06)\gamma^{2}$ MPa, shown in Fig. 11. This fit has a coefficient of determination of 0.99, corresponding to a standard deviation of 0.5 MPa around the fit, indicating that the quadratic scaling is a good approximation. The coefficient for the model curvature is consistent with the fit value within the 37% model uncertainty. Thus, we can attribute the stress-strain relation of packings of staples to the deformation of particles due to tensile loadings.
For completeness, we tested a quadratic fit with two fit parameters which yields $a_{1}=(2.71\pm 0.04)$ MPa and $a_{2}=(59.4\pm 0.2)$ MPa. In this case, the coefficient $a_{2}$ only changes by a few percent, and $a_{1}$ only contributes to $9\%$ of the total stress at $\gamma=0.5$. Thus, the strength of packings of entangling staples can also be approximated by a dominant quadratric function without a linear term, which is again attributed to a probability of entanglement $p_{s}$ proportional to increasing strain $\gamma$, similar to chains. With the linear term, the data fits with a reduced $\chi^{2}=1.0$ when using the standard deviation of multiple runs as an uncertainty, indicating the linear term is reproducible in the case of staples, if relatively small.
@@ VI Discussion
It is interesting to consider if this understanding of entanglement could lead to mixtures with enhanced strength or strain stiffening in soils, concrete, and other granular materials. Fibers are typically only added to soils at a fraction less than about 1% to provide reinforcement, and while this reinforcement increases strength, it does not produce strain-stiffening when the fibers are mixed in randomly et al (2022). The small fraction of fibers is likely too little to produce strong strain stiffening from system-filling clusters of entanglements. The same could be said about concrete reinforced with fibers AHFW92. Larger packing fractions of fibers have been confirmed to enhance the weaker self-amplified friction mechanism for strain stiffening from localized strain Dumont et al. (2018).
Another model of for athermal fibrous gels produces strong strain stiffening Sharma et al. (2016). That system is modeled as an arrangement of particles with elastic bonds with tension and bending stiffness between some fraction of neighbors. Under strain, rearrangements are mostly prevented by the bonds, so increasing strain on the system increases the number of stretched bonds, resulting in strain stiffening. The source of stress is analogous to our case, where the increasing stretching results in strain stiffening, although in our case it is physical entanglements that prevents rearrangement and cause more entangled particles to be deformed instead of gel bonds. The fiber model requires a numerical simulation, in contrast to our model which breaks down the stress contributions into a low-dimensional model using Eq. 2. In the fiber model, the fraction of bonds with neighbors is assumed to be enough that the system would be above the gel percolation threshold, which would correspond to greater than one connection per node in the Erdos-Renyi model. This parameter space is mathematically analogous to our experiments with system filling clusters of entanglements ($N>8$). The fiber model emphasizes that the strain stiffening is a critical behavior with a critical transition in strain, which is different from the critical transition we find with increasing $N$. Our system does not have such a critical transition in strain, as the contribution of the elastic membrane to the stress-strain curve is dominant at small strains. More generally, in other macroscopic particle granular systems the small-strain region is usually dominated by a confining stress which is required to prevent the system from collapse due to its self-weight (even highly entangled systems tend to have a partial collapse without some confinement).
@@ VII Conclusions
Using x-ray tomography measurements of the packing structure, we showed that the number of entanglements $n$ of chains is nearly proportional to an effective free enclosed area $A_{eff}$ surrounded by a chain available to entangle other particles, accounting for excluded volume effects (Fig. 6). This area $A_{eff}$ and the expected number of entanglements $n$ can be calculated given the shape of the particles in the packing assuming random particle arrangements (Fig. 5). The threshold of about 1 entanglement per enclosure required to produce system-filling clusters and strong strain stiffening is similar to the Erdös-Rényi model for random graphs, which is often used to describe the gelation transition.
We presented this model in a form that could be extended to packings of different particle shapes that can entangle. In particular, for a particular rigid particle geometry, one could calculate from a Monte Carlo simulation the possible entanglement arrangements the effective free enclosed area $A_{eff}$ (Appendix E), which leads to the number of entanglements $n$, and thus whether or not there are enough entanglements to connect in system-filling clusters to produce strong strain stiffening. While these quantities can be predicted from the known geometry of rigid particles (Fig. 15), calculating these quantities required detailed structural measurements for flexible chains to calculate $A_{eff}$ and $n$.
@@
We presented a model for the stress-strain relation of random packings with system-filling entanglements, where the stress originates from the stress required to deform the particles, which is due to tensional loads on the particles in the cases tested. Using a mean-field model, the slope of the stress-strain curve of the packing was written as a product of an effective particle modulus, the packing fraction, the probability of stretched particles $p_{s}$, and the ratio of particle strain to packing strain (Eq. 2). For both chains and staples, the increase in effective stiffness with strain that characterizes strain stiffening can be attributed to the increase in the fraction of stretched particles $p_{s}(\gamma)$ as packing strain $\gamma$ increases (Fig. 7b, 13) and causes more entangled particles to pull on each other, while the effective particle stiffness and the ratio of particle strain to packing strain are approximately constant (Figs. 17, 24, 14). A model with a linear increase in the probability of stretched particles $p_{s}$ with packing strain $\gamma$ corresponds to a quadratic stress-strain relation, which is a decent approximation of the observed stress-strain curves. The curvature of the stress-strain relation was predicted within the model uncertainties ($\leq 37\%$) of measured values for packings of both chains and staples (Figs. 3, 11). To predict the stress-strain relation, the effective particle modulus and packing fraction are relatively easy to measure, however, the probability of stretching and the ratio of particle strain to packing strain required information from experiments with three-dimensional imaging, permanent plastic deformation, or simulations of packings, and are not yet easily predictable from first principles.
@@ VIII Acknowledgements
We thank Mike Dunlap for assistance with materials testing equipment, Peter Eng and Mark Rivers for their assistance with x-ray tomography at the Advanced Photon Source, Ling-Nan Zhou for his particle tracking code, Nicholas Rodenberg and Dylan Murphy for preliminary measurements, Sulimon Sattari for performing simulations of network models for comparison, Henos Musie for performing measurements for Fig. 24, Tammer Abiyu for performing measurements for Fig. 11, David Brantley for performing measurements for Fig. 18, and Xuan-Truc Nguyen for analyzing particles for Figs. 12 and 14. We acknowledge the Advanced Photon Source general user program at Argonne National Laboratory for providing beam time. This work was supported by the NSF MRSEC program under DMR-0820054 and DMR-2011854.
@@ Appendix A Self-amplified friction model comparison
Why is the scaling of our stress-strain curves so different from those of Dumont et al. (2018)? An important difference is that Dumont et al. (2018) used a narrow indenter with an open top surface that allows dilation of the packing without resistance, which results in a more localized strain so a few chains near the indenter are responsible for most of the displacement. We confirmed experimentally that using a narrow indenter and free boundary allows beads to move past each other when the packing is indented. This more localized motion allows the mechanism of self-amplified friction to determine which chains move and thus the response measured by the indenter. In contrast, our experiments used a flat plate that produced a relatively uniform strain in the packing, and a weak confinement from the sides with the elastic membrane. In this case, the entanglements of long chains prevent particles from shearing past each other, thus preventing the self-amplified friction mechanism, and instead requiring load to be generated by deformation of particles. This comparison confirms the importance of preventing shear banding Regev and Reichhardt (2013) in the strong strain stiffening behavior that we observed. However, it also highlights that the apparent measured stress-strain relation is not necessarily an intrinsic material property, and even different mechanisms can exist for strain stiffening for the same material depending on whether the boundary conditions allow a more localized shear banding or prevents it.
@@ Appendix C Bend angles of chains
To help determine whether the compression of tight loops of chains contributes to strong strain stiffening, the bending stiffness of a chain loop was measured by pulling on a looped chain. A chain with $\xi=8$ beads was bent into a loop (shown in the inset of Fig. 18) and held in place between two sheets of acrylic separated by 4.5 mm in such a manner that the chain would not unloop or catch on another bead as the ends of the chain were pulled. Pictures were taken to measure local bend angles $\theta$ between the links of neighboring chains (shown in the inset of Fig. 18) with an uncertainty of 1 degree limited by the image resolution. Four angles were measured for each chain (those centered on the beads with solid circles overlaid in their centers in the inset of Fig. 18), and the measurement was repeated for two different chains to obtain a total of eight measured angles at several different extensional force values. The average angles for each force are shown in Fig. 18, with the error bars indicating the standard deviations of the measured angles. We measured no resolvable force until the chains were bent beyond $45\pm 1^{\circ}$ degrees, which corresponds to the minimum loop circumference $\xi=360^{\circ}/45^{\circ}=8.0$. While there is an initial linear force $F_{\theta}$ with bend angle $\theta$, above 2.5 N, the chains started to bend non-uniformly, indicated by the large increase in standard deviation of $\theta$ with increasing force shown in Fig. 18; in particular one link would bend much more than the others, indicating yielding of the weakest link.
@@ Appendix D Packing fraction
The global packing fraction was also obtained from the x-ray tomography data at strains $\gamma=0$ and $0.2$ by dividing the sphere volume by the average volume of convex hulls surrounding the beads. The $N=5$ and $N=20$ samples dilated ($\phi$ decreased) under compression, while the $N=10$ sample was initially loose (presumably due to an inconsistency in packing procedure) so compacted ($\phi$ increased) under compression. The packing fraction at strains of $\gamma=0.2$ appears to follow a smooth trend with $N$, which is consistent with the expectation that packings evolve towards a critical packing fraction (a.k.a. critical state) under shear, regardless of initial packing fraction (Scholfield and Wroth, 1968). The initially loose packing for $N=10$ may have contributed to there being very few links at full extension at $\gamma=0$ (Fig. 7b), which made it impossible to resolve the average link strain $\gamma_{p}$ for $N=10$, and about half the number of entanglements per enclosure as other datasets at $\gamma=0$ (Fig. 6). Once the sample was strained to $\gamma=0.2$, these statistics matched more closely with the data with the other strain stiffening packing at $N=20$.
@@ Appendix E Calculation of effective free enclosed area $A_{eff}$
The ratio of $A_{eff}/A$ is plotted as a function of chain segment length $m$ in Fig. 23. We find a consistent $A_{eff}/A=0.27\pm 0.01$ in the range of strong strain stiffening ($m\geq 8$), with the $\pm$ corresponding to the standard deviation at different contour lengths $m$ for both $N=10$ and 20. This consistency of the ratio $A_{eff}/A$ in this range is in agreement with the earlier expectation that this ratio is a constant for staples Gravish et al. (2012) ($A_{eff}/A$ corresponds to $1/\alpha$ in Gravish et al. (2012)). In the limit of large $m$, the distance from where the entangled chain crosses the entangling manifold to the entangling particle $l(\Phi,x)$ is likely to be large compared to the particle diameter $a$. In that limit, the factor $1-a^{2}/l(\Phi,x)^{2}$ in the integrand of Eq. 10 approaches 1. For chains, this corresponds to the limit of large $m$ which can have entanglement manifolds much wider than $a$. However, for smaller $m\leq 8$, the ratio $A_{eff}/A$ does drop to zero as the constraints of excluded volume become more significant where the entangled width of the entanglement manifold is only slightly larger than $a$ [i.e. the largest $l(\Phi,x)$ is only slightly larger than $a$]. Using the derivation of Eq. 10 as a guide, it can be inferred that the maximum possible value of $A_{eff}/A$ is 1/2 in the extreme limit of thin entangled particles inside a large entangled volume where $a/l(\Phi,x)$ would be small for most of the volume.
@@ E.1 Monte Carlo calculations of $A_{eff}/A$
The calculation of $A_{eff}/A$ for chains shows that the number of entanglements $n$ can be predicted given a particle geometry, but since chains have a flexible shape it required input of x-ray tomography data to get the particle geometry in the packing. To confirm that such a geometric calculation for $A_{eff}/A$ can make a priori predictions for rigid particles based on the particle shape without packing information, we carried out Monte Carlo simulations of $A_{eff}/A$ for staples at various aspect ratios of arm length over backbone length. The geometry of staples is different than Fig. 22, assuming cylindrical arms of diameter $a$ in the enclosed rectangular area $A=(L_{x}-a/2)(L_{y}-a/2)$. We randomly sample values of arms crossing the entangled plane at positions $20\leq x\leq L_{x}-a/$ 2, $20\leq y\leq L_{y}/2-a/$ 2, and orientations $20\leq\Theta\leq\pi/$ 2 (with probability weighted by $\sin\Theta$), and $20\leq\Phi\leq\pi/$ 2. Valid entanglements with no overlap of particles in the set of coordinates are weighted by a factor of $\cos\Theta$ in the average to account for the larger volume of the staple in the entangling volume, and contribute a factor of zero if there is overlap. A set of positions and orientations is considered to be a valid entangled position without overlapping another particle if
@@
The values of $A_{eff}/A$ obtained from this Monte Carlo calculation for staples are shown in Fig. 15 for different aspect ratios $L_{x}/L_{y}$ with a fixed backbone length $L_{y}=13a$, corresponding to the simulation of Gravish et al. (2012).